# In Search of Truth: An Interrogation Approach to Hallucination Detection

### v2

@@ A.2 More results for variable $K$ values
Tab. 9 and 10 depicts the performance of InterrogateLLM on the books and GCI datasets, with variable $K$ values. As can be seen, higher values of $K$ result in enhanced accuracy.
@@ A.3 More variable temperature results
Tab. 6 and 7 depicts the performance of the InterrogateLLM method with and without variable temperature values.
@@ B.3 GCI Dataset Query-Answer Examples
Table 6: Results for Books dataset with same temperature and variable temperature.
@@
Table 7: Results for GCI dataset with same temperature and variable temperature.
@@
Table 8: Results for the Movies dataset. Results are reported for different k values, ranging from 1 to 5, average and max scores.
@@
Table 9: Results for the Books dataset. Results are reported for different k values, ranging from 1 to 5, with average and max scoring.
@@
Table 10: Results for the GCI dataset. Results are reported for different k values, ranging between 1 to 5, average scoring.

### v3

@@ A.2 More results for variable $K$ values
Tab. 8, 9 and 10. depicts the performance of InterrogateLLM on the books and GCI datasets, with variable $K$ values. As can be seen, higher values of $K$ result in enhanced accuracy.
@@ A.3 More variable temperature results
Tab. 5, 6 and 7 depicts the performance of the InterrogateLLM method with and without variable temperature values.
@@ B.3 GCI Dataset Query-Answer Examples
Table 5: Results for Movies dataset, presenting results for constant and variable temperature, with both average and maximum scores.
@@
Table 6: Results for Books dataset, presenting results for constant and variable temperature, with both average and maximum scores.
@@
Table 7: Results for GCI dataset, presenting results for constant and variable temperature, with both average and maximum scores.
@@
Table 8: Evaluation results for the Movies dataset across different k values (1 to 5), with average and maximum scores presented.
@@
Table 9: Evaluation results for the Books dataset across different k values (1 to 5), with average and maximum scores presented.
@@
Table 10: Evaluation results for the GCI dataset across different k values (1 to 5), with average and maximum scores presented.